OpenStudy (anonymous):

What number must you add to complete the square? x2 + 14x = -7 A. 196 B. 49 C. 7 D. 14

myininaya (myininaya):

\[x^2+bx+(b/2)^2=(x+b/2)^2\]

myininaya (myininaya):

so add to both sides \[(b/2)^2 \] can you guess what b should be?

OpenStudy (anonymous):

14x ?

myininaya (myininaya):

well b is just the 14 bx is 14x but our formula calls for just b which is 14

myininaya (myininaya):

so plug 14 into that (b/2)^2 thing to figure out your answer as to what to add to both sides in order to complete the square

OpenStudy (anonymous):

oh 196

myininaya (myininaya):

no

myininaya (myininaya):

\[(\frac{14}{2})^2 \] what is 14/2?

OpenStudy (anonymous):

7 so the answer would be 49

myininaya (myininaya):

\[(\frac{14}{2})^2=(7)^2=49\] right